# Analysis of Linear Mode Connectivity via Permutation-Based Weight Matching: With Insights into Other Permutation Search Methods

### v4

@@ Appendix B Calculation of $R$ with Threshold $\gamma$
In Section 4.4, we explained that even when the $L_{2}$ distance between the weights of two models is large, their outputs can be close depending on the input distribution. Here, we use a simple example for a more detailed analysis.
@@ Appendix D Experimental Setup
This section describes the experimental setup for training neural networks to obtain SGD solutions. We apply Sinkhorn’s algorithm for permutation based on WM and STE. Thus, we also provide detailed information on the experimental setup for Sinkhorn’s algorithm. Four datasets were used in this study: MNIST (Lecun et al., 1998), Fashion-MNIST (FMNIST) (Xiao et al., 2017), CIFAR10 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009).
All experiments were conducted on a Linux workstation with two AMD EPYC 7543 32-Core processors, eight NVIDIA A30 GPUs, and 512 GB of memory. The PyTorch 2.1.0222https://pytorch.org/, PyTorch Lightning 2.1.0333https://lightning.ai/docs/pytorch/stable/, and torchvision 0.16.0444https://pytorch.org/vision/stable/index.html libraries were used for model training and evaluation.
@@ D.1 Model Training
We utilized the VGG16 and ResNet20 architectures of (Ainsworth et al., 2023). To accomplish Layer-wise Model Compression (LMC), we increased the widths of VGG11 and ResNet20 by factors of 4 and 16, respectively. As described in (Jordan et al., 2023), we used the training dataset to repair the BatchNorm layers in these models during model merging. Optimization was conducted using Adam with a learning rate of $1\times 10^{-3}$. The batch size and maximum number of epochs were set to 512 and 100, respectively. The following data augmentations were performed during training: random $32\times 32$ pixel crops, and random horizontal flips.
@@ D.2 Permutation Search
For permutation search in WM and STE, we employed the method based on Sinkhorn’s algorithm as proposed by Peña et al. (2023). We utilized the implementation provided by the authors in their GitHub repository [^6] https://github.com/fagp/sinkhorn-rebasin. DistL2Loss and MidLoss were used as loss functions for permutation searches corresponding to WM and STE, respectively. Optimization was performed using Adam with a learning rate of 1 for MLP, VGG11, and ResNet20 and 10 for ResNet50, setting the maximum number of epochs to 10 for DistL2Loss and five for MidLoss. For MidLoss, the batch size was set to 512; for DistL2Loss, there was no batch size because the dataset was not used. 100 iterations of parameter updates were performed per epoch for DistL2Loss.
@@ F.2 Matrix Representation of Convolutional Layer
There exists a matrix $\bm{M}$ such that $\mathrm{vec}(\bm{Y})=\bm{M}\mathrm{vec}(\bm{X})$ holds (Sedghi et al., 2019; Jain, 1989; Goodfellow et al., 2016), where
@@ Appendix G Proofs
Consider the $L_{2}$ norm of the $\ell$-th layer $\|\bm{W}_{\ell}^{(a)}-\bm{P}_{\ell}\bm{W}_{\ell}^{(b)}\bm{P}_{\ell-1}^{\top}\|^{2}$. Using the fact that the $L_{2}$ norm can be rewritten using trace, we have
@@
In convolutional layers, permutation matrices permute the input and output channels of the kernel. Therefore, the permutation matrices $\bm{P}_{\ell}$ and $\bm{P}_{\ell-1}$ corresponding to the input and output are $m\times m$ matrices. By using these matrices, the permutation of the matrix representation of the convolutional layer of the model $\bm{\theta}_{b}$ is denoted by $(\bm{P}_{\ell}\otimes\bm{I}_{m})\bm{M}^{(b)}(\bm{P}_{\ell-1}\otimes\bm{I}_{m})^{\top}$. Lemma G.1 indicates that finding the permutation matrices that minimize the $L_{2}$ distance between the two kernels is equivalent to minimizing $\|\bm{M}^{(a)}-(\bm{P}_{\ell}\otimes\bm{I}_{m})\bm{M}^{(b)}(\bm{P}_{\ell-1}\otimes\bm{I}_{m})^{\top}\|$. Therefore, we have
@@
holds, where we let $\bm{C}_{w}^{(a)}=\sum_{i}s_{w,i}^{(a)}\bm{u}_{w,i}^{(a)}(\bm{v}_{w,i}^{(a)})^{*}$ and $\bm{C}_{w}^{(b)}=\sum_{i}s_{w,i}^{(b)}(\bm{P}_{\ell}\bm{u}_{w,i}^{(a)})(\bm{P}_{\ell-1}\bm{v}_{w,i}^{(a)})^{*}$. From Equation 59 and Equation 62, we have
@@
Using the fact that if $w\neq w^{\prime}$, then $\bm{e}_{w}^{\top}\bm{e}_{w^{\prime}}=0$, and otherwise, $\bm{e}_{w}^{\top}\bm{e}_{w^{\prime}}=1$, we have
@@ Appendix H Additional Experimental results
H.1 Distribution of Singular Values
@@
Figure 7 shows the distributions of the singular values of all the layers. The figure demonstrates that, in all layers except for the output layer, the singular values are very similar across all models. Meanwhile, in the output layer, there is variability in the singular values. However, this variability does not affect the accuracy of the merged model. Let $\bm{W}_{L}^{(a)}=\sum_{i}s^{(a)}_{L,i}\bm{u}^{(a)}_{L,i}(\bm{v}^{(a)})_{L,i}^{\top}$ and $\bm{W}_{L}^{(b)}=\sum_{i}s^{(b)}_{L,i}\bm{u}^{(b)}_{L,i}(\bm{v}^{(b)}_{L,i})^{\top}$ represent the output layer weights of the two trained models. The figure shows that the difference between the singular values of the two models is approximately a constant multiple. In other words, there exists a constant $\alpha$ such that $s_{L,i}^{(a)}\approx\alpha s_{L,i}^{(b)}$ for all $i$. If the singular vectors of the two weights are equal (i.e., $\bm{v}^{(a)}_{L,i}=\bm{v}^{(b)}_{L,i}$ and $\bm{u}^{(a)}_{L,i}=\bm{u}^{(b)}_{L,i}$ for all $i$), then $\bm{W}_{L}^{(a)}\approx\alpha\bm{W}_{L}^{(b)}$ holds (indeed, as mentioned in Section 4.3, the permutation matrix aligns the directions of the singular vectors). Therefore, the weight of the output layer of the merged model at the ratio $\lambda\in[0,1]$ is given by $\lambda\bm{W}_{L}^{(a)}+(1-\lambda)\bm{W}_{L}^{(b)}\approx\lambda\bm{W}_{L}^{(a)}+(1-\lambda)\alpha\bm{W}_{L}^{(a)}=(\lambda+(1-\lambda)\alpha)\bm{W}_{L}^{(a)}$. Thus, we can consider that the weight and the activation function of the merged model are given by $\bm{W}_{L}^{(a)}$ and a softmax function with an inverse temperature of $1/(\lambda+(1-\lambda)\alpha)$, respectively. Since the inverse temperature does not affect the accuracy value, the difference in the singular values of the output layer would not matter in satisfying LMC, at least in terms of accuracy.
H.2 Inner Products Between Right Singular Vectors of Hidden Layers And Their Input
@@
H.3 Relationship with Model Width
@@
We also investigated the effect of model width on the inner products between the hidden layer inputs and the right singular vectors. Figures 10 and 11 show the values of these inner products for each layer as model width changes. Figures 10 and 11 show the distributions of inner products for VGG11 and ResNet20 models trained on CIFAR10, respectively. These figures demonstrate that as model width increases, the inner products between the right singular vectors with large singular values and the inputs also increase.
@@
First, focusing on the results in Figure 12 with $\gamma=0$, we observe that the value of $R$ decreases even when the width increases and WM is used. Conversely, Figure 12 shows that the directions of singular vectors with particularly large singular values are aligned by permutation as model width increases. This suggests that even with WM, it is difficult to perfectly align the directions of singular vectors between the two models. However, increasing the width decreases the fraction of singular vectors with large singular values, and thus, making it easier for WM to align the directions of these dominant singular vectors.
As shown in Figure 9, when the model is sufficiently wide, the proportion of large singular values is very small compared to the total number of singular values. Furthermore, Figures 10 and 11 demonstrate that the right singular vectors associated with these relatively large singular values have a large inner product with the hidden layer input. This means that the number of singular vectors that WM needs to align to achieve LMC is reduced when the model is wide enough, as discussed in Section 4.4. Indeed, Figure 12 suggests that WM preferentially aligns these significant singular vectors. Therefore, increasing the width is expected to make LMC more feasible.
H.4 Dependency of Weight Decay and Learning Rate
@@
In Section 4, we stated that the reason why WM can establish LMC is that the permutation found by WM aligns singular vectors with large singular values between two models. In other words, the smaller the proportion of large singular values in the weights of each layer of the models, the more likely LMC is to be established by WM. In fact, some previous studies (Galanti et al., 2023; Timor et al., 2023) have shown that increasing the weight decay and learning rate during model training can reduce the ranks of the weights in the trained model. Therefore, it is highly likely that the results observed by Qu & Horvath (2024) were caused by the reduction in the ranks of the weights. In the following, we will experimentally confirm this prediction.
@@
From Figure 13, we can see that the ratio of large singular values decreases as both the weight decay and the learning rate increase. This has already been noted in previous research (Galanti et al., 2023; Timor et al., 2023). Additionally, Figure 13 shows that the test loss of the merged model decreases when both the test loss of the original model and the ratio of large singular values are small. This can be explained by our analysis in Section 4. As we described, WM facilitates LMC by aligning the singular vectors between the two models and making the functions of the middle layers of the merged model and the original model more similar. In other words, this suggests that it is challenging for the merged model to outperform the original model. Furthermore, the smaller the ratio of large singular values, the easier it is to align the singular vectors using WM, so the functions of the merged model and the original model become more closely aligned. From this, we conclude that the higher the performance of the original model and the smaller the ratio of large singular values, the better the performance of the merged model. This is consistent with the results shown in Figure 13.
H.5 Activation Matching
@@ H.6 STE and WM
In this subsection, additional experimental results for Section 6.3 are shown in Table 4 for the barrier values and $R$ values between each pair of models. The table shows the model-merging results with $\lambda=1/2$, and the mean and standard deviation of five model merges. In the table, a negative value for the barrier indicates an improvement in performance due to the merging. From the Table 4, we can see that the barrier between $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$ is also smaller for WM than for STE.
H.7 Dependency of $R$ on Threshold $\gamma$
@@
H.8 LMC on ResNet50 trained on ImageNet
@@
(b) $L_{2}$ distance between $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$.
$L_{2}$ dist. w/ WM $L_{2}$ dist. w/o WM
@@
Table 5 presents the results of merging ResNet50 models trained on the ImageNet dataset. Table 5 shows the test loss and top-1 accuracy of the models before and after merging (i.e., the pre-merged models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, and the merged model $\bm{\theta}_{c}$). Table 5 shows the $L_{2}$ distance between models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ before and after applying permutations. These tables report the mean and standard deviation for five independent model merges. According to Table 5, the test loss and accuracy of the merged model are clearly improved by using WM. However, they are still worse than those of the pre-merged models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, indicating that the LMC cannot be considered satisfied. Table 5 demonstrates that using WM decreases the $L_{2}$ distance between models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$. This decrease in $L_{2}$ distance is larger than that observed for VGG11 and ResNet20 in Table 1, suggesting that the singular vectors of models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ are better aligned by permutations. Figure 17 presents the results of evaluating $R$ for each model pair. When $\gamma=0.3$, Figure 17 shows that $R(\bm{\theta}_{a},\bm{\theta}_{b})$ rises to about 0.5 with WM, and the value of $R$ increases to approximately 0.9 between the pre- and post-merged models (i.e., $R(\bm{\theta}_{a},\bm{\theta}_{c})$ and $R(\bm{\theta}_{b},\bm{\theta}_{c})$). Thus, even with ResNet50, the singular vectors with large singular values are aligned between the pre- and post-merged models by using WM.
@@
To investigate why LMC does not hold even though the singular vectors with large singular values are aligned by using WM, we examine the distributions of singular values at each layer and the inner products between right singular vectors and the inputs. Figure 18 shows the distribution of singular values for each layer, and Figure 19 shows the distribution of the average absolute values of the inner products between the right singular vectors and the inputs for each layer. Each figure is plotted in a different color for the 10 trained models. Figure 18 demonstrates that the distributions of singular values are nearly identical across all models. Thus, the difference in singular values between models is not a reason for preventing LMC. In Figure 19, focusing on the distribution of the inner product between the right singular vectors and the inputs, we observe that the right singular vectors with large singular values does not necessarily have a large inner product with the input. As discussed in Section 4.4, WM can only align singular vectors with large singular values, so this discrepancy can be a reason preventing the establishment of LMC. As shown in Figures 11 and 10, the wider the model, the larger the inner product of the input and the right singular vectors with large singular values in the hidden layers. Therefore, it is considered necessary to increase the width of the model to establish LMC with ResNet50.

### v5

@@ Appendix B Calculation of $R$ with Threshold $\gamma$
In Section 4.4, we explained that even when the $L^{2}$ distance between the weights of two models is large, their outputs can be close depending on the input distribution. Here, we use a simple example for a more detailed analysis.
@@ Appendix D Experimental Setup
This section describes the experimental setup for training neural networks to obtain SGD solutions. We apply Sinkhorn’s algorithm for permutation based on WM and STE. Thus, we also provide detailed information on the experimental setup for Sinkhorn’s algorithm. Four datasets were used in this study: MNIST (Lecun et al., 1998), Fashion-MNIST (FMNIST) (Xiao et al., 2017), CIFAR10 (Krizhevsky & Hinton, 2009), and ImageNet (Deng et al., 2009).
All experiments were conducted on a Linux workstation with two AMD EPYC 7543 32-Core processors, eight NVIDIA A30 GPUs, and 512 GB of memory. The PyTorch 2.1.0444https://pytorch.org/, PyTorch Lightning 2.1.0555https://lightning.ai/docs/pytorch/stable/, and torchvision 0.16.0666https://pytorch.org/vision/stable/index.html libraries were used for model training and evaluation.
@@ D.1 Model Training
We utilized the VGG16 and ResNet20 architectures of (Ainsworth et al., 2023). To accomplish Linear Mode Connectivity (LMC), we increased the widths of VGG11 and ResNet20 by factors of 4 and 16, respectively. As described in (Jordan et al., 2023), we used the training dataset to repair the BatchNorm layers in these models during model merging. Optimization was conducted using Adam with a learning rate of $1\times 10^{-3}$. The batch size and maximum number of epochs were set to 512 and 100, respectively. The following data augmentations were performed during training: random $32\times 32$ pixel crops, and random horizontal flips.
@@ D.2 Permutation Search
For permutation search in WM and STE, we employed the method based on Sinkhorn’s algorithm as proposed by Guerrero-Peña et al. (2023). We utilized the implementation provided by the authors in their GitHub repository [^8] https://github.com/fagp/sinkhorn-rebasin. DistL2Loss and MidLoss were used as loss functions for permutation searches corresponding to WM and STE, respectively. Optimization was performed using Adam with a learning rate of 1 for MLP, VGG11, and ResNet20 and 10 for ResNet50, setting the maximum number of epochs to 10 for DistL2Loss and five for MidLoss. For MidLoss, the batch size was set to 512; for DistL2Loss, there was no batch size because the dataset was not used. 100 iterations of parameter updates were performed per epoch for DistL2Loss.
@@ F.2 Matrix Representation of Convolutional Layer
There exists a matrix $\bm{M}$ such that $\mathrm{vec}(\bm{Y})=\bm{M}\mathrm{vec}(\bm{X})$ holds (Sedghi et al., 2019; Jain, 1989; Salakhutdinov, 2014), where
@@ Appendix G Proofs
Consider the $L^{2}$ norm of the $\ell$-th layer $\|\bm{W}_{\ell}^{(a)}-\bm{P}_{\ell}\bm{W}_{\ell}^{(b)}\bm{P}_{\ell-1}^{\top}\|^{2}$. Using the fact that the $L^{2}$ norm can be rewritten using trace, we have
@@
In convolutional layers, permutation matrices permute the input and output channels of the kernel. Therefore, the permutation matrices $\bm{P}_{\ell}$ and $\bm{P}_{\ell-1}$ corresponding to the input and output are $m\times m$ matrices. By using these matrices, the permutation of the matrix representation of the convolutional layer of the model $\bm{\theta}_{b}$ is denoted by $(\bm{P}_{\ell}\otimes\bm{I}_{m})\bm{M}^{(b)}(\bm{P}_{\ell-1}\otimes\bm{I}_{m})^{\top}$. Lemma G.1 indicates that finding the permutation matrices that minimize the $L^{2}$ distance between the two kernels is equivalent to minimizing $\|\bm{M}^{(a)}-(\bm{P}_{\ell}\otimes\bm{I}_{m})\bm{M}^{(b)}(\bm{P}_{\ell-1}\otimes\bm{I}_{m})^{\top}\|$. Therefore, we have
@@
hold, where we let $\bm{C}_{w}^{(a)}=\sum_{i}s_{w,i}^{(a)}\bm{u}_{w,i}^{(a)}(\bm{v}_{w,i}^{(a)})^{*}$ and $\bm{C}_{w}^{(b)}=\sum_{i}s_{w,i}^{(b)}(\bm{P}_{\ell}\bm{u}_{w,i}^{(a)})(\bm{P}_{\ell-1}\bm{v}_{w,i}^{(a)})^{*}$. From Equation 59 and Equation 62, we have
@@
Using the fact that if $w\neq w^{\prime}$, then $\bm{e}_{w}^{\top}\bm{e}_{w^{\prime}}=0$, and otherwise, $\bm{e}_{w}^{\top}\bm{e}_{w^{\prime}}=1$, we get
@@ Appendix H Additional Experimental results
H.1 Learning curve of WM
In this subsection, Figure 6 shows the learning curves for VGG11 and ResNet20 when WM is performed using Sinkhorn’s algorithm. The figure shows that the distance between the two models decreases as the training progresses, and the performance of the merged model also improves. In this paper, for both VGG11 and ResNet20, we used permutations at the 10th epoch, when the loss of the merged model is stably small.
(a) Results of VGG11
(b) Results of ResNet20
Figure 6: $L^{2}$ distance between two models, test loss, and accuracy of merged models when optimizing permutations using Sinkhorn’s algorithm for WM. Five permutation search trials were conducted with independently trained models (i.e., ten independently trained models were prepared to form five model pairs, and WM was performed for each pair). These results are plotted in different colors.
H.2 Distribution of Singular Values
@@
Figure 8: Adjusted distributions of the singular values of the output layer.
Figure 7 shows the distributions of the singular values of all the layers. The figure demonstrates that, in all layers except for the output layer, the singular values are very similar across all models. Meanwhile, in the output layer, there is variability in the singular values. However, this variability does not affect the accuracy of the merged model. Let $\bm{W}_{L}^{(a)}=\sum_{i}s^{(a)}_{L,i}\bm{u}^{(a)}_{L,i}(\bm{v}^{(a)})_{L,i}^{\top}$ and $\bm{W}_{L}^{(b)}=\sum_{i}s^{(b)}_{L,i}\bm{u}^{(b)}_{L,i}(\bm{v}^{(b)}_{L,i})^{\top}$ represent the output layer weights of the two trained models. The figure shows that the difference between the singular values of the two models is approximately a constant multiple. In other words, there exists a constant $\alpha$ such that $s_{L,i}^{(a)}\approx\alpha s_{L,i}^{(b)}$ for all $i$. To confirm this, Figure 8 shows the distribution of singular values when the constant $\alpha$ is calculated and the weight of the output layer is adjusted, demonstrating that correcting the output layer by a constant factor can address the differences in the distribution. If the singular vectors of the two weights are equal (i.e., $\bm{v}^{(a)}_{L,i}=\bm{v}^{(b)}_{L,i}$ and $\bm{u}^{(a)}_{L,i}=\bm{u}^{(b)}_{L,i}$ for all $i$), then $\bm{W}_{L}^{(a)}\approx\alpha\bm{W}_{L}^{(b)}$ holds (indeed, as mentioned in Section 4.3, the permutation matrix aligns the directions of the singular vectors). Therefore, the weight of the output layer of the merged model at the ratio $\lambda\in[0,1]$ is given by $\lambda\bm{W}_{L}^{(a)}+(1-\lambda)\bm{W}_{L}^{(b)}\approx\lambda\bm{W}_{L}^{(a)}+(1-\lambda)\alpha\bm{W}_{L}^{(a)}=(\lambda+(1-\lambda)\alpha)\bm{W}_{L}^{(a)}$. Thus, we can consider that the weight and the activation function of the merged model are given by $\bm{W}_{L}^{(a)}$ and a softmax function with an inverse temperature of $1/(\lambda+(1-\lambda)\alpha)$, respectively. Since the inverse temperature does not affect the accuracy value, the difference in the singular values of the output layer would not matter in satisfying LMC, at least in terms of accuracy.
H.3 Inner Products Between Right Singular Vectors of Hidden Layers And Their Input
@@
H.4 Relationship with Model Width
@@
We also investigated the effect of model width on the inner products between the hidden layer inputs and the right singular vectors. Figures 11 and 12 show the values of these inner products for each layer as model width changes. Figures 11 and 12 show the distributions of inner products for VGG11 and ResNet20 models trained on CIFAR10, respectively. These figures demonstrate that as model width increases, the inner products between the right singular vectors with large singular values and the inputs also increase.
@@
First, focusing on the results in Figure 13 with $\gamma=0$, we observe that the value of $R$ decreases even when the width increases and WM is used. Conversely, Figure 13 shows that the directions of singular vectors with particularly large singular values are aligned by permutation as model width increases. This suggests that even with WM, it is difficult to perfectly align the directions of singular vectors between the two models. However, increasing the width decreases the fraction of singular vectors with large singular values, thus making it easier for WM to align the directions of these dominant singular vectors.
As shown in Figure 10, when the model is sufficiently wide, the proportion of large singular values is very small compared to the total number of singular values. Furthermore, Figures 11 and 12 demonstrate that the right singular vectors associated with these relatively large singular values have a large inner product with the hidden layer input. This means that the number of singular vectors that WM needs to align to achieve LMC is reduced when the model is wide enough, as discussed in Section 4.4. Indeed, Figure 13 suggests that WM preferentially aligns these significant singular vectors. Therefore, increasing the width is expected to make LMC more feasible.
H.5 Dependency of Weight Decay and Learning Rate
@@
In Section 4, we stated that the reason why WM can establish LMC is that the permutation found by WM aligns singular vectors with large singular values between two models. In other words, the smaller the proportion of large singular values in the weights of each layer of the models, the more likely LMC is to be established by WM. In fact, some previous studies (Galanti et al., 2022; Timor et al., 2023) have shown that increasing the weight decay and learning rate during model training can reduce the ranks of the weights in the trained model. Therefore, it is highly likely that the results observed by Qu & Horvath (2024) were caused by the reduction in the ranks of the weights. In the following, we will experimentally confirm this prediction.
@@
From Figure 14, we can see that the ratio of large singular values decreases as both the weight decay and the learning rate increase. This has already been noted in previous research (Galanti et al., 2022; Timor et al., 2023). Additionally, Figure 14 shows that the test loss of the merged model decreases when both the test loss of the original model and the ratio of large singular values are small. This can be explained by our analysis in Section 4. As we described, WM facilitates LMC by aligning the singular vectors between the two models and making the functions of the middle layers of the merged model and the original model more similar. In other words, this suggests that it is challenging for the merged model to outperform the original model. Furthermore, the smaller the ratio of large singular values, the easier it is to align the singular vectors using WM, so the functions of the merged model and the original model become more closely aligned. From this, we conclude that the higher the performance of the original model and the smaller the ratio of large singular values, the better the performance of the merged model. This is consistent with the results shown in Figure 14.
H.6 Activation Matching
@@ H.7 STE and WM
(a) Results of WM.
(b) Results of STE.
In this subsection, additional experimental results for Section 6.3 are shown in Table 5 for the barrier values between each pair of models. The table shows the model-merging results with $\lambda=1/2$, and the mean and standard deviation of three model merges. In the table, a negative value for the barrier indicates an improvement in performance due to the merging. In addition, Figure 17 shows the accuracy landscape around $\bm{\theta}_{a}$, $\pi_{b}(\bm{\theta}_{b})$, and $\pi_{c}(\bm{\theta}_{c})$. From Tables 5 and 17, we can see that the barrier between $\pi_{b}(\bm{\theta}_{b})$ and $\pi_{c}(\bm{\theta}_{c})$ is also smaller for WM than for STE.
H.8 Dependency of $R$ on Threshold $\gamma$
@@
H.9 LMC on ResNet50 trained on ImageNet
@@
(b) $L^{2}$ distance between $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$.
$L^{2}$ dist. w/ WM $L^{2}$ dist. w/o WM
@@
Table 6 presents the results of merging ResNet50 models trained on the ImageNet dataset. Table 6 shows the test loss and top-1 accuracy of the models before and after merging (i.e., the pre-merged models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, and the merged model $\bm{\theta}_{c}$). Table 6 shows the $L^{2}$ distance between models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ before and after applying permutations. These tables report the mean and standard deviation for five independent model merges. According to Table 6, the test loss and accuracy of the merged model are clearly improved by using WM. However, they are still worse than those of the pre-merged models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, indicating that the LMC cannot be considered satisfied. Table 6 demonstrates that using WM decreases the $L^{2}$ distance between models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$. This decrease in $L^{2}$ distance is larger than that observed for VGG11 and ResNet20 in Table 1, suggesting that the singular vectors of models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ are better aligned by permutations. Figure 19 presents the results of evaluating $R$ for each model pair. When $\gamma=0.3$, Figure 19 shows that $R(\bm{\theta}_{a},\bm{\theta}_{b})$ rises to about 0.5 with WM, and the value of $R$ increases to approximately 0.9 between the pre- and post-merged models (i.e., $R(\bm{\theta}_{a},\bm{\theta}_{c})$ and $R(\bm{\theta}_{b},\bm{\theta}_{c})$). Thus, even with ResNet50, the singular vectors with large singular values are aligned between the pre- and post-merged models by using WM.
@@
To investigate why LMC does not hold even though the singular vectors with large singular values are aligned by using WM, we examine the distributions of singular values at each layer and the inner products between the right singular vectors and the inputs. Figure 20 shows the distribution of singular values for each layer, and Figure 21 shows the distribution of the average absolute values of the inner products between the right singular vectors and the inputs for each layer. Each figure is plotted in a different color for the 10 trained models. Figure 20 demonstrates that the distributions of singular values are nearly identical across all models. Thus, the difference in singular values between models is not a reason for preventing LMC. In Figure 21, focusing on the distribution of the inner product between the right singular vectors and the inputs, we observe that the right singular vectors with large singular values do not necessarily have a large inner product with the input. As discussed in Section 4.4, WM can only align singular vectors with large singular values, so this discrepancy can be a reason preventing the establishment of LMC. As shown in Figures 12 and 11, the wider the model, the larger the inner product of the input and the right singular vectors with large singular values in the hidden layers. Therefore, it is considered necessary to increase the width of the model to establish LMC with ResNet50.